# Samson Lasaulce, Giuseppe Valenzise, Zhuo He, and Merouane Debbah

### v1

@@ VI Challenges
VI-A Prior information combination
One critical challenge lies in the effective utilization of prior information to enhance the accuracy of generated samples from sparse observations. In RF sensing, additional prior information, such as spatial configurations, historical signal patterns, and environmental characteristics, is often available and can significantly improve reconstruction accuracy by providing context and reducing ambiguity in the generated data. To leverage this, it is essential to integrate the prior information seamlessly with the architecture design in generative models, such as GANs or diffusion models, possibly by conditioning the generative model based on the prior information.
VI-B Distortion-perception tradeoff
GenAI techniques are highly effective in tasks requiring human or machine perception but encounter challenges when distortion is more critical, such as in radio map reconstruction. Therefore one challenge is to develop models that optimize both distortion and perception. One possible way to address this challenge is to integrate traditional neural networks, trained to minimize distortion, with GenAI models when dealing with problems that involve both distortion and perception.
VI-C Implementation of large models
Another challenge is the storage and computing of large GenAI models, particularly LLMs. These models enable intelligent edges in IoT systems but present storage and computing challenges on user equipment or edge devices. Efficient model compression techniques, such as pruning, quantization, or more developed techniques such as goal-oriented compression, are essential to deploy large models in resource-constrained equipment. Overcoming this challenge is essential for the practical deployment of GenAI-empowered RF sensing techniques.
VI-D Balancing synthetic and real data
While GenAI offers significant potential to enhance the training of neural networks by generating large amounts of synthetic data, it is crucial to carefully consider the balance between synthetic and real-world data. A major concern in the context of data augmentation with GenAI is determining the optimal amount of synthetic data to use in training a neural network. An insufficient amount of synthetic data can lead to inadequate training and poor generalization, while an excessive amount of synthetic data may introduce biases and reduce the model’s performance on real-world data.

### v2

@@ VI Challenges
VI-A Model Deployment
Deploying generative models in RF sensing is challenging due to high computational demands and the limited resources of IoT devices and wireless sensor networks. Models like Transformers and diffusion models often exceed the processing power and memory of lightweight RF hardware, raising feasibility concerns [14]. Real-time processing complicates matters further by requiring low-latency responses. Solutions include edge computing and task offloading, where edge servers or cloud nodes manage intensive tasks, allowing IoT devices to focus on lighter functions [15]. Model compression techniques like quantization and pruning can also reduce model size and computational needs, making deployment on constrained devices more feasible. Additionally, using lightweight, task-specific models for functions such as localization or anomaly detection minimizes device load and enhances adaptability. These strategies collectively help balance performance with the constraints of resource-limited environments.
VI-B Model Interpretability
Generative models’ complex architectures result in ”black box” characteristics, limiting interpretability and applicability in high-stakes domains like autonomous navigation. Simplification via distillation or pruning, along with probabilistic outputs to represent prediction confidence, can improve interpretability. Additionally, generative outputs may lack physical validity within RF contexts. Incorporating physics-informed constraints, such as propagation principles, into model structures or loss functions can ensure results align with real-world RF behaviors, enhancing robustness and reliability.
VI-C Data Requirements
Generative AI models in RF sensing demand large, diverse datasets to capture varied environmental conditions, barrier types, and signal fluctuations across IoT settings. Data collection can be labor-intensive, especially when labeled data is scarce or conditions are dynamic, as seen in smart cities and industrial sites. Sparse or incomplete RF data further limits model generalization. To address this, transfer learning, few-shot training, and data augmentation techniques—such as simulation-based methods—can synthetically enrich training datasets, reducing dependency on extensive real-world data.
VI-D Prior information combination
A key challenge is effectively utilizing prior information to enhance the accuracy of generated samples from sparse observations. In RF sensing, additional prior information—such as spatial configurations, historical signal patterns, and environmental characteristics—can significantly improve reconstruction accuracy by providing context and reducing ambiguity. To leverage this, it is crucial to integrate prior information seamlessly into the architecture design of generative models. By employing the inherent statistical properties of certain Generative AI models, explainable techniques with statistical frameworks, like physics-informed neural networks, can enhance performance while requiring fewer training resources.